Shift report 20000817 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Thu Aug 17 2000 - 23:44:52 EDT

  • Next message: http account: "Shift report 20000818 00:00-08:00"

    ------------------------------------------------------------
    Date: Thu Aug 17 23:44:52 2000, Shift: 16:00-24:00
    Supervisor: RW, with: KO,B3
    Log book#4 pages 112 - 114
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam was cogged at 19:55 with I_b=9.6e9 and I_y=10.9e9.
       Trigger 1 rate was around 60. The beam is still going on.
    
    2. Note that D1 and D2 polarity was reversed with the same
       magnitude before this shift.
    
       RD wanted to optimize the anode voltage of MTCP1.
       Different voltage was applied on MTCP1.
    
       MTCP1       anode(V)   no of events
       run 2325     1100V       20K (trigger 1)
       run 2326     1150V       20K
       run 2327     1250V       20K
       run 2329     1300V       37K    fb1 taken out
       run 2330     1275V       20K
       run 2332     1225V       20K
       run 2333     1200V       20K
       run 2334     1175V       20K
       run 2335     1125V       20K
    
       Survey data run finished. Anode voltage is set 1200V
       the same value as that before this survey.
    
       run 2336  started at 11:37 with the trigger 1 rate of
       20-25.
    
    3. Dead time was found 75%. KO found that fb1 caused the 
       problem. fb1 is taken out from the DAQ (It is doing 
       nothing, according to KO). Dead time becomes about 
       20-25% at the trigger rate of 100 (75-80 events are 
       taken).
    
       T3 cannot hold any voltage. Wire might be broken. 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Thu Aug 17 2000 - 23:48:28 EDT